OpenStudy (anonymous):

Explain. 3x + 7 = 1/2X + 2

OpenStudy (anonymous):

get all like variables to one side, and then divide to get X alone. subtract 1/2x to get rid of it on the right side, then subtract 7 to get ride of it on the left side. then divide both sides by 2 1/2 to get x= 3x + 7 = 1/2X + 2 -1/2x -1/2x 2 1/2x + 7 = 2 -7 -7 2 1/2x = -5 divide both sides by 2 1/2 x = -2